Staff:

LBTO: Juan Carlos

Observer: Shark-Nir team (local at the telescope)

Summary.

Continuing the pre-com3 for shark-nr.

Acquisitons of NCPA and coro centering algorithm.

There were two events of PSF moved away, and the loop was not open. Those events have to be checked with the telemetry (To be checked the offload.)

Argos cal unit, deployment ok the park positions , the park switch is not activated.

Shark-log:

https://docs.google.com/document/d/1fYgeR-pu_lNgZE4gLrYJ3335OtDDleGAxxuGYEzEOm0/edit#heading=h.xvpj69e9dfvn

Details.

7:30 AM

Telescope released for Shark-Nir daytime.

7:40am

Power on WFS.

PDU: 5,8,15,15 ON

hexpod positions: 0.8,0.4,-1.15,356,-674

fov.conf #8.

OSS didn't move the swing arm.

move using the argos interface.

8:40am

PresetAO
  • Loop close without disturbance.
  • gain:1.75,0.5,0.4
  • gopton
  • less vibraitons today
Shark confirm good PSF to work with the PD.

bayside stage movemebts at close loop.
  • new:
  • x->65.7
  • y:->20.95
  • z=na
shark taking data and adjustment of the PSF in teh detector.
  • offset in Z for phase diversity.
  • initial: z0=62
  • z1=59

Disturbance on.
  • s2.0
  • optmize gain:20221119_165055
  • 5.38,1.08,0.78
  • gopton
  • modalSR: 85
modalpot_s2_lbti_20221119.png

move z to 59

move to 62

10:50am

Z move to 59

11:08

Z move to 62

11:37

Z move to 59

TN:20221119_184335

Z move to 62
  • Shark PSF
  • Filter H2
  • pscale:0.0145arcsec

shark_psf_PD.png

11:51 am

TEst coronographic

TN:20221119_185258

:_193132

_193153

10K frames.

moving motors in shark.

check for jitter introduced by shark (?).

TN:_194118

_194405

stop coro procedure.

Shark team analysing procedures, etc..

starting PD

_194705

_194734

_195013

Z stage 59

1:02PM

Z move to 62

_200238

_200417

Move stage from: X:65.7 to 65.9, Y:20.95 to 20.85

Mirror ventilation started at: ~12:30 ( to be checked.)

_201207

loop open.

Light gone.

(?)

PresetAO
  • Optimize:202651
  • 6.22,1.08,0.78
  • freq 130Hz.>10mas
  • gopt on
  • _203048

repeat optmize:203103

optmize_lbti_b1_s1_203103.png
  • 6.36,1.08,0.63
  • visible 130Hz.
  • 203452
  • gopt gain:2.48,0.42,0.27
  • one click down TT
  • 2.38TT (better)
  • _203627
  • _203811
  • _200238
01:30PM

Start coro alignment.

Effect of mirror ventilation, ncrease the tilt by 2.

jitter_mirrorventilation_on_20221119.png

residual modal plot.

modalplot_20221119_0151.png

PSf moved away.
  • loop open
  • lopp closed
  • No disturbance
  • manual optimze: 1.85,0.45,0.35.
TN:212854
  • to check all the offloads.
Adjust the bayside stage.
  • move bay side to align the PSF: 66.3, 21.0
  • _214109
move stage:
  • 65.9,20.85
  • _215001
2:55pm

pupil alignment

Stop 3:20PM

Power off PDU:

5,8,15,16. Left on 11 (bayside stage).

Argos swing arm parked. encoder 56.85, but switch not activated.

-- JuanCarlosGuerra - 19 Nov 2022
Topic revision: r3 - 20 Nov 2022, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback